# Quantum-Coherent Regime of Programmable Dipolar Spin Ice

### 30 Mar 2026 v1

@@ Abstract
The term spin ice describes magnetic systems with an extensively degenerate ground-state manifold, characterized by a finite residual entropy analogous to that of water ice [Pauling1928]. The minimum energy configurations obey local ice rules, most notably the “two-in–two-out” constraint at each vertex, like the four hydrogen bonds, with two protons close and two distant from each oxygen atom. These enforce a divergence-free condition on an emergent field. Violations of this constraint generate mobile, fractionalized excitations that carry effective magnetic charge and interact via Coulomb-like potentials [Castelnovo2011].
@@
Artificial spin ices (ASIs) are engineered realizations of such constrained systems, originally inspired by frustrated rare-earth pyrochlores [Ramirez1999, Bramwell2001]. These are usually two dimensional systems, including a square ASI based on the square lattice of point-like dipoles, obeying the “two-in–two-out” ice rule [Lieb1967], and a Kagome ASI, which is based on a hexagonal lattice [Gartside2018]. Traditionally, artificial spin ice has been implemented using lithographically patterned nanomagnets [Wang2006, tanaka2006magnetic, Farhan2019, Parakkat2021, Goryca2021, Jensen2024], or 3D nanowire device [May2021], as well as platforms based on superconducting vortices [Wang2018], colloids [Han2008, Rodriguez-Gallo2023], and other classical realizations [Miao2020MonopoleGas, Meeussen2020].These systems have enabled direct imaging of monopoles and Dirac strings, but experimental access to their quantum-coherent dynamics has remained elusive.
@@
We build upon recent developments, which enabled the implementation of spin ice physics in reconfigurable quantum annealers, where the dipoles are represented by superconducting qubits [King2021, Zhou2021]. This approach allows not only precise control of geometry and coupling strengths but also the ability to pin individual spins, strongly impose Gauss-law boundary conditions, and introduce topological charges at will. Unlike classical nano-magnetic realizations, the system is inherently subject to both thermal and quantum fluctuations, which can activate monopole dynamics and drive the system across its low-energy manifold, characterized by the monopole charge conservation. However, these implementations have only been able to explore classical artificial spin ices and relied on strictly short-range, vertex-level couplings, omitting the dipolar interactions that play a central role in natural and other artificial spin ices. Dipolar interactions lift degeneracies, stabilize ordered phases, and modify monopole kinetics through long-range energetic and entropic interaction potentials [Nisoli2013]. Incorporating these interactions is therefore essential for approaching the full phenomenology of dipolar spin ice and, ultimately, regimes related to quantum spin-liquid physics.
Leveraging the latest generation of superconducting-qubit quantum annealers, we realize a programmable dipolar spin-ice model, with a direct mapping between lattice spins and physical qubits and engineer effective dipolar interactions. This allows us to study the real-time dynamics, activated by fluctuations originating from tunable transverse-field, of Dirac-string defects and monopole plasmas on frustrated lattices exceeding 400 vertices. We access a previously unexplored quantum-coherent regime of artificial spin ice. Tracking the probability distributions of motion of individual defects, we observe super-diffusive monopole transport, with scaling exponents intermediate between classical diffusion and ballistic motion. This behaviour signals dynamics beyond classical stochastic relaxation and indicates coherent propagation of fractionalized defects within a constrained gauge manifold. We further investigate the collective evolution of many interacting defects, realizing a dynamical magnetic monopole plasma whose expansion and correlations reflect the interplay of gauge constraints and dipolar interactions. Together, these results establish programmable quantum spin ice on superconducting annealers as a new experimental platform, enabling direct studies of fractionalized excitations, entropic Coulomb interactions and emergent gauge fields in engineered quantum many-body systems. Such phenomena may ultimately find applications in electronics [Heyderman2013] or low-energy computation [Hon2021, Edwards2023, Caravelli2022], with quantum collective effects making these prospects even more compelling.
@@
Figure 1: Schematic of the artificial spin ice model and its ground states. a: Each yellow dot represents a position of single point dipole (qubit). Straight black arrows indicate the dipole orientations. In subsequent panels dipoles that are flipped compared to this configuration are highlighted in yellow. Four types of interaction terms are included, representatives are pictured as curved arrows. Interactions values are set to ones corresponding to truncation of two‑dimensional dipolar interactions to the first four terms. (b-d): Ice‑rule ground states obtained from annealing for different model parameters. Holes in the lattice correspond to device defects. In all ground states, each vertex satisfies the ice rule, indicated by two yellow and two black colored arrows. Vertex statistics and correlations extracted from annealing are compared with the six‑vertex ice ground states (c), in order to identify the point $J_{\perp}=J_{\parallel}$. In the following, we focus on the configuration shown in (b) and investigate the dynamics of its excitations.
@@ Methods
We study an ASI Hamiltonian inspired by a square lattice of point-like dipoles. The geometry, conventions, and interaction terms defining the Hamiltonian are shown in Fig. 1a. Each dipole is mapped onto a single flux qubit of the D‑Wave Advantage2 quantum annealer. This mapping describe in detail in Supplementary Material is possible due to increased connectivity of newer device [Advantage2], over its predecessors. The quantum annealing device consists of a network of superconducting flux qubits that implements the transverse‑field Ising Model, represente by the Hamiltonian:
@@
where $\mathcal{J}(t)$ and $\Gamma(t)$ are time‑dependent schedules controlling, respectively, the strength of the classical Ising and transverse-field contributions, $\langle ij\rangle$ indicates a sum over interacting pairs of spins $i$ and $j$. $\hat{\sigma}_{i}^{z}$ and $\hat{\sigma}_{i}^{x}$ are the Pauli operators acting on spin $i$ in the $z$ and $x$ directions respectively. $J_{ij}$ and $h_{i}$ denote the programmable couplers and local bias fields, whose values remain fixed during the evolution scaled only by the global $\mathcal{J}(t)$ and $\Gamma(t)$. Couplings $J_{ij}$ are implemented only between physically connected qubits, corresponding to overlapping or adjacent superconducting loops. These qubits are fabricated as elongated superconducting flux loops arranged in two orthogonal layers [Bunyk2014], which enables efficient embedding of two-dimensional lattice models onto the hardware connectivity graph [Boo2021]. The longitudinal($z$) Ising energy scale $\mathcal{J}(t)$ and the transverse‑field($x$) strength $\Gamma(t)$ are driven globally by a time dependent waveform. The local connections: $J_{\perp}^{\prime}=J_{\perp}+C$ and $J_{\parallel}^{\prime}=J_{\parallel}+C$ consist of a constant offset $C$, which enforces charge conservation and constrains the dynamics to the target Hilbert space, together with additional terms arising from the truncated expansion of the two‑dimensional dipolar interactions.
@@
Previous implementations of artificial spin-ice models on quantum annealers encoded each effective dipole using strongly coupled chains of qubits  [King2021, Zhou2021]. In these older method, a dipole flip requires a collective reversal of an entire qubit chain, constituting a higher‑order tunnelling process. As a consequence, quantum fluctuations were strongly suppressed, preventing direct observation of intrinsic system dynamics. In contrast, our approach encodes each dipole in a single qubit, such that a dipole flip maps directly onto a single‑qubit transition driven by the local transverse field, enabling direct access to the system’s dynamics.
@@
The terms J,J $_{\perp},J_{\parallel}\text{ and }J_{4}$ are implemented via the couplers available on the device. The interaction $J_{3}$ is implemented indirectly: the corresponding dipole-dipole coupling arises from the time‑averaged fast dynamics of an intermediary qubit that effectively mediates the otherwise missing connection. This approach constitutes an alternative strategy for enhancing connectivity relative to previous qubit chain method and is particularly advantageous here, as it avoids the higher‑order tunneling that would otherwise suppress the system dynamics. Technical details of $J_{3}$ implementation are provided in the Supplementary Material. The terms $J_{3}$ and $J_{4}$ are extensions of the usual fully local framework. The device permits these terms to have different signs from those typically available for $J_{\perp},J_{\parallel}$ providing a new route to engineer more complex interacting artificial spin liquids. This capability enables access to regimes that are inaccessible in nano-magnetic systems or in previous annealing experiments. In the present work, we restrict attention interections strength corresponding to a two‑dimensional dipolar interactions truncated to these four coupling types.
@@
We realized a lattice comprising of 23x23 vertices, with 28 vertices removed due to physical device defects that prevented implementation of the local Hamiltonian. The measurements were run on the Advantage2 system version 1.5. Since then, device performance has since degraded, with additional qubits becoming inoperable with the current version being 1.13. We provide the full embedding code for mapping the square‑ice model onto the Zephyr connectivity graph of the quantum annealer, together with all scripts required to reproduce the experiments, in a publicly accessible digital archive [^1] Giergiel, Krzysztof (2026): Quantum-Coherent Regime of Programmable Dipolar Spin Ice. CSIRO. v1. Software. https://doi.org/10.25919/vraf-2p08.
@@
The previous artificial-ice platforms have demonstrated magnetic monopoles[Miao2020MonopoleGas], Dirac strings[Morris2009, Mengotti2011], and classical Coulomb phases[Farhan2019], but have lacked the programmability or dynamical control needed to tune long-range interactions or explore Z2 spin-liquid behavior, particularly in the non-classical regime. We are able to prepare desired initial states, quickly increase quantum fluctuations to a desired strength and for a prescribed amount of time, and again freeze the dynamics for the readout. The states are prepared and read out in the Z basis, with the pair-wise ZZ interactions of the model Hamiltonian, and the X field supplying the fluctuations. Repeated measurements allow reconstruction of the probability distribution generated by the fluctuation driven dynamics. A schematic overview of the experimental protocol is provided in Fig. 2.
@@
To tune the parameters, we first used standard forward annealing, which generated a sample of ground states. Based on the sample we computed the vertex type statistics. The expected vertex statistics depend on the ratio of $J_{\perp}/J_{\parallel}$ [Coraux2024] and the local bias fields [Goryca2021]. Representative vertex types and ice configurations obtained from annealing across different parameter regimes are shown in Fig. 1. The parameters were tuned to reproduce vertex statistics consistent with pure ice‑rule behavior ($J=0$) in the absence of bias fields, corresponding to the highly degenerate six‑vertex spin‑ice model. These parameters values correspond to the condition we describe as non-interacting.
We chose to keep the timing of the experimental sequence fixed, using the fastest available slopes to reach the desired level of fluctuations. The system was then held at the desired level of fluctuations for $\Delta t=100~\text{ns}$. We modify only the strength of fluctuations $\Gamma$, while proportionally adjusting the ZZ terms implementing the dipolar interactions terms. This is equivalent to tuning an effective evolution time $t_{\mathrm{eff}}=\Gamma\Delta t$. The conversion to physical units is approximately $1~\text{(a.u.)}\approx 0.124$, corresponding to $0.124~\text{GHz}\times 100~\text{ns}$ at $100~\text{(a.u.)}$. It takes about $250~\text{ns}$ to reach fluctuation corresponding to $t_{eff}=100$. The ZZ interactions enforcing the ice rules were fixed at 45% of total available range of the device, at the point where we observed the lowest generation and recombination rates of the monopole charges. The scaled ZZ terms implementing dipolar interactions were added on top of this baseline. In all experiments, the system dynamics were initiated from the line‑ordered state (Fig. 1b), to which the desired defects were introduced (see Fig. 2).
Figure 2: Schematic illustration of measurement sequence. The black colors on plots here are plotted, when the dipole state is opposite to the ground state configuration (Fig. 1b). Circles correspond to the vertices with $\pm 2Q$ charge (3-1 in-out or 1-3 in-out). Vertices with charges $\pm 4Q$ were very suppressed. Initially the system dynamics are frozen by high values of interaction terms. We change the parameters as fast as possible decreasing the $ZZ$ interaction terms and increasing the $\Gamma X$ - fluctuation terms. The system is kept fixed $\Delta t=t_{2}-t_{1}=100~\text{ns}$ in the evolving condition. We vary the fluctuations strength $\Gamma$ to obtain different effective interactions time $t_{\text{eff}}=\Gamma\Delta t$. The $J_{\perp},J_{\parallel},J_{3},J_{4}$ parameters are rescaled accordingly. The ice rules are always enforced at $45\%$ of total available value of ZZZZ interactions. For example for $t_{\text{eff}}=100$ the energy associated with generation of pair of charges is still large $E_{2Q}=20\Gamma$, but for longer effective times ice rules are enforced less strictly.
@@ Results
Both exponents are smaller than the ballistic value of $\alpha=2$ expected for coherent free propagation, yet exceed than the classical diffusive value of $\alpha=1$. Moreover, the significant difference between single-particle displacement and relative separation indicates correlated motion between the two monopoles. This behavior emerges in a regime where explicit dipolar couplings are set to zero and only the ice-rule constraint is enforced. In a purely classical dynamics respecting the same constraint, would correspond to thermally activated random walks along allowed spin-flip paths, yielding diffusive scaling. The observed super-diffusive but sub-ballistic transport therefore indicates dynamics beyond classical stochastic relaxation.
@@
Additional illustrations in the online only Appendix Fig. A1 and A2 extend this analysis to system with different charge densities and varying interaction conditions. The condition of plasma is similar to popping a balloon filled with plasma in a vacuum. Increasing monopole plasma density or decreasing dipolar interactions causes the cloud to expand faster. The additional effect of dipolar interactions is evident in increased short-range correlations between positive and negative charges.
Qualitative comparison with Z2 lattice gauge theory shows that the observed screening behavior, correlation decay, and absence of long confining strings align with expectations for a deconfined Z2 spin liquid. We provide a first observation of microscopic dynamics of quantum spin-liquid of spin-ice defects, but we do not provide a full understanding of the particular values of transport exponents. Anomalous dynamics observed even in minimally interacting regimes arise from irreducible entropic interactions imposed by the constrained Hilbert space.
@@
[…]
Supplementary information Appended at the end of this file:
Programmable Dipolar Quantum Spin Ice on a Superconducting Qubit Annealer: Supplementary Information
@@
[…]
Declarations
  • Funding
    The Polish National Science Centre (NCN) Sonata Bis grant 2019/34/E/ST3/00405. Access to the D-Wave system was funded by the CSIRO Quantum Technologies Future Science Platform.
  • Conflict of interest/Competing interests
    The authors declare no competing interests.
  • Ethics approval and consent to participate
    Not applicable
  • Consent for publication
    Not applicable

### 19 May 2026 v2

@@ Abstract
The term spin ice describes magnetic systems with an extensively degenerate ground-state manifold, characterized by a finite residual entropy analogous to that of water ice [29]. The minimum energy configurations obey local ice rules, most notably the “two-in–two-out” constraint at each vertex, like the four hydrogen bonds, with two protons close and two distant from each oxygen atom. These enforce a divergence-free condition on an emergent field. Violations of this constraint generate mobile, fractionalized excitations that carry effective magnetic charge and interact via Coulomb-like potentials [6].
@@
Artificial spin ices (ASIs) are engineered realizations of such constrained systems, originally inspired by frustrated rare-earth pyrochlores [31, 2]. These are usually two dimensional systems, including a square ASI based on the square lattice of point-like dipoles, obeying the “two-in–two-out” ice rule [19], and a Kagome ASI, which is based on a hexagonal lattice [12]. Traditionally, artificial spin ice has been implemented using lithographically patterned nanomagnets [34, 33, 11, 28, 13, 17], or 3D nanowire device [20], as well as platforms based on superconducting vortices [35], colloids [14, 32], and other classical realizations [24, 22].These systems have enabled direct imaging of monopoles and Dirac strings, but experimental access to their quantum-coherent dynamics has remained elusive.
@@
We build upon recent developments, which enabled the implementation of spin ice physics in reconfigurable quantum annealers, where the dipoles are represented by superconducting qubits [18, 36]. This approach allows for not only precise control of geometry and coupling strengths, but also the ability to pin individual spins, strongly impose Gauss-law boundary conditions, and introduce topological charges at will. Unlike classical nano-magnetic realizations, the system is inherently subject to both thermal and quantum fluctuations, which can activate monopole dynamics and drive the system across its low-energy manifold, characterized by the monopole charge conservation. However, these implementations have only been able to explore classical artificial spin ices and have relied on strictly short-range, vertex-level couplings, omitting the dipolar interactions that play a central role in natural and other artificial spin ices. Dipolar interactions lift degeneracies, stabilize ordered phases, and modify monopole kinetics through long-range energetic and entropic interaction potentials [26]. Incorporating these interactions is therefore essential for approaching the full phenomenology of dipolar spin ice and, ultimately, regimes related to quantum spin-liquid physics.
@@
Figure 1: Schematic of the artificial spin ice model and its ground states. a: Each yellow dot represents a position of single point dipole (qubit). Straight black arrows indicate the dipole orientations. In subsequent panels dipoles that are flipped compared to this configuration are highlighted in yellow. Four types of interaction terms are included, representatives are pictured as curved arrows. (b-d): Ice‑rule ground states obtained from annealing for different model parameters.
Leveraging the latest generation of superconducting-qubit quantum annealers, we realize a programmable dipolar spin-ice model, with a direct mapping between lattice spins and physical qubits and engineer effective dipolar interactions. This allows us to study the real-time dynamics, activated by fluctuations originating from tunable transverse field, of Dirac-string defects and monopole plasmas on frustrated lattices exceeding 400 vertices. We access a previously unexplored quantum-coherent regime of artificial spin ice. Tracking the probability distributions of motion of individual defects, we observe super-diffusive monopole transport, with THE scaling exponents intermediate between classical diffusion and ballistic motion. This behavior signals dynamics beyond classical stochastic relaxation and indicates coherent propagation of fractionalized defects within a constrained gauge manifold. We further investigate the collective evolution of many interacting defects, realizing a dynamical magnetic monopole plasma whose expansion and correlations reflect the interplay of gauge constraints and dipolar interactions. Together, these results establish programmable quantum spin ice on superconducting annealers as a new experimental platform, enabling direct studies of fractionalized excitations, entropic Coulomb interactions and emergent gauge fields in engineered quantum many-body systems. Such phenomena may ultimately find applications in electronics [15] or low-energy computation [16, 10, 4], with quantum collective effects making these prospects even more compelling.
@@ Methods
Figure 2: Schematic illustration of measurement sequence. The black colors on plots here are plotted, when the dipole state is opposite to the ground state configuration (Fig. 1b). Circles correspond to the vertices with $\pm 2Q$ charge (3-1 in-out or 1-3 in-out). Vertices with charges $\pm 4Q$ were very suppressed. Initially the system dynamics are frozen by high values of interaction terms. We change the parameters as fast as possible decreasing the $ZZ$ interaction terms and increasing the $\Gamma X$ - fluctuation terms. The system is kept fixed $\Delta t=t_{2}-t_{1}=100\penalty 10000\ \text{ns}$ in the evolving condition. We vary the fluctuations strength $\Gamma$ to obtain different effective interactions time $t_{\text{eff}}=\Gamma\Delta t$. The $J_{\perp},J_{\parallel},J_{3},J_{4}$ parameters are rescaled accordingly. The ice rules are always enforced at $45\%$ of total available value of ZZZZ interactions. For example for $t_{\text{eff}}=100$ the energy associated with generation of pair of charges is still large $E_{2Q}=20\Gamma$, but for longer effective times ice rules are enforced less strictly.
We study an ASI Hamiltonian inspired by a square lattice of point-like dipoles. The geometry, conventions, and interaction terms defining the Hamiltonian are shown in Fig. 1a. Each dipole is mapped onto a single flux qubit of the D‑Wave Advantage2 quantum annealer. This mapping, described in detail in the Supplementary Material, is possible due to the increased connectivity of the newer device [21], over its predecessors. The quantum annealing device consists of a network of superconducting flux qubits that implements the transverse‑field Ising Model, represented by the Hamiltonian:
@@
where $\mathcal{J}(t)$ and $\Gamma(t)$ are time‑dependent schedules controlling, respectively, the strength of the classical Ising and transverse-field contributions, $\langle ij\rangle$ indicates a sum over interacting pairs of spins $i$ and $j$. $\hat{\sigma}_{i}^{z}$ and $\hat{\sigma}_{i}^{x}$ are the Pauli operators acting on spin $i$ in the $z$ and $x$ directions respectively. $J_{ij}$ and $h_{i}$ denote the programmable couplers and local bias fields, whose values remain fixed during the evolution scaled only by the global $\mathcal{J}(t)$ and $\Gamma(t)$. Couplings $J_{ij}$ are implemented only between physically connected qubits, corresponding to overlapping or adjacent superconducting loops. These qubits are fabricated as elongated superconducting flux loops arranged in two orthogonal layers [3], which enables efficient embedding of two-dimensional lattice models onto the hardware connectivity graph [1]. The longitudinal($z$) Ising energy scale $\mathcal{J}(t)$ and the transverse‑field($x$) strength $\Gamma(t)$ are driven globally by a time dependent waveform. The local connections: $J_{\perp}^{\prime}=J_{\perp}+C$ and $J_{\parallel}^{\prime}=J_{\parallel}+C$ consist of a constant offset $C$, which enforces charge conservation and constrains the dynamics to the target Hilbert space, together with additional terms arising from the truncated expansion of the two‑dimensional dipolar interactions.
@@
Previous implementations of artificial spin-ice models on quantum annealers encoded each effective dipole using strongly coupled chains of qubits [18, 36]. In this older method, a dipole flip requires a collective reversal of an entire qubit chain, constituting a higher‑order tunneling process. As a consequence, quantum fluctuations were strongly suppressed, preventing direct observation of intrinsic system dynamics. In contrast, our approach encodes each dipole in a single qubit, such that a dipole flip maps directly onto a single‑qubit transition driven by the local transverse field, enabling direct access to the system’s dynamics.
@@
The terms J,J $_{\perp},J_{\parallel}\text{ and }J_{4}$ are implemented via the couplers available on the device. The interaction $J_{3}$ is implemented indirectly: the corresponding dipole-dipole coupling arises from the time‑averaged fast dynamics of an intermediary qubit that effectively mediates the otherwise missing connection. This approach constitutes an alternative strategy for enhancing connectivity relative to the previous qubit chain method and is particularly advantageous here, as it avoids the higher‑order tunneling that would otherwise suppress the system dynamics. Technical details of $J_{3}$ implementation are provided in the Supplementary Material. The terms $J_{3}$ and $J_{4}$ are extensions of the usual fully local framework. The device permits these terms to have different signs from those typically available for $J_{\perp},J_{\parallel}$ providing a new route to engineer more complex interacting artificial spin liquids. This capability enables access to regimes that are inaccessible in nano-magnetic systems or in previous annealing experiments. In the present work, we restrict attention to a single choice of interaction strengths corresponding to a two-dimensional dipolar interaction truncated to these four coupling types. For dipolar interaction strength $U$ the total couplings are: $J_{\perp}^{\prime}=C+2U,\penalty 10000\ J_{\parallel}^{\prime}=C+U,\penalty 10000\ J_{3}=-U,\penalty 10000\ J_{4}=-\frac{1}{4}U$.
@@
We realized a lattice comprising of 23x23 vertices, with 28 vertices removed due to physical device defects that prevented implementation of the local Hamiltonian. The measurements were run on the Advantage2 system version 1.5. Since then, device performance has since degraded, with additional qubits becoming inoperable with the current version being 1.13. We provide the complete embedding code for mapping the square‑ice model onto the Zephyr connectivity graph of the quantum annealer, together with all scripts required to reproduce the experiments, in a publicly accessible digital archive [^1] Giergiel, Krzysztof (2026): Quantum-Coherent Regime of Programmable Dipolar Spin Ice. CSIRO. v1. Software. https://doi.org/10.25919/vraf-2p08.
@@
The previous artificial-ice platforms have demonstrated magnetic monopoles [24], Dirac strings [25, 23], and classical Coulomb phases [11], but have lacked the programmability or dynamical control needed to tune long-range interactions or explore Z2 spin-liquid behavior, particularly in the non-classical regime. We are able to prepare desired initial states, quickly increase quantum fluctuations to a desired strength and for a prescribed amount of time, and again freeze the dynamics for the readout. The states are prepared and read out in the Z basis, with the pair-wise ZZ interactions of the model Hamiltonian, and the X field supplying the fluctuations. Repeated measurements allow reconstruction of the probability distribution generated by the fluctuation driven dynamics. A schematic overview of the experimental protocol is provided in Fig. 2.
@@
To tune the parameters, we first used standard forward annealing, which generated a sample of ground states. Based on the sample we computed the vertex type statistics. The expected vertex statistics depend on the ratio of $J_{\perp}/J_{\parallel}$ [8] and the local bias fields [13]. Representative vertex types and ice configurations obtained from annealing across different parameter regimes are shown in Fig. 1. The parameters were tuned to reproduce vertex statistics consistent with pure ice‑rule behavior ($J=0$) in the absence of bias fields, corresponding to the highly degenerate six‑vertex spin‑ice model. These parameters values correspond to the condition we describe as non-interacting.
We chose to keep the timing of the experimental sequence fixed, using the fastest available slopes to reach the desired level of fluctuations. The system was then held at the desired level of fluctuations for $\Delta t=100\penalty 10000\ \text{ns}$. We modify only the strength of fluctuations $\Gamma$, while proportionally adjusting the ZZ terms implementing the dipolar interactions terms. This is equivalent to tuning an effective evolution time $t_{\mathrm{eff}}=\Gamma\Delta t$. The conversion to physical units is approximately $1\penalty 10000\ \text{(a.u.)}\approx 0.124$, corresponding to $0.124\penalty 10000\ \text{GHz}\times 100\penalty 10000\ \text{ns}$ at $100\penalty 10000\ \text{(a.u.)}$. It takes about $250\penalty 10000\ \text{ns}$ to reach fluctuation corresponding to $t_{eff}=100$. The ZZ interactions enforcing the ice rules were fixed at 45% of total available range of the device, at the point where we observed the lowest generation and recombination rates of the monopole charges. The scaled ZZ terms implementing dipolar interactions were added on top of this baseline. In all experiments, the system dynamics were initiated from the line‑ordered state (Fig. 1b), to which the desired defects were introduced (see Fig. 2).
@@ Results
Both exponents are smaller than the ballistic value of $\alpha=2$ expected for coherent free propagation, yet exceed than the classical diffusive value of $\alpha=1$. Moreover, the significant difference between single-particle displacement and relative separation indicates correlated motion between the two monopoles. This behavior emerges in a regime where explicit dipolar couplings are set to zero and only the ice-rule constraint is enforced.
In a purely classical dynamics respecting the same constraint, monopole motion would correspond to thermally activated random walks along allowed spin-flip paths. Such dynamics are expected to be initially sub-diffusive, with an exponent $\alpha\approx 0.86$ based on numerical simulations [7], close to $\alpha\approx 0.85$ predicted by simple Glauber spin dynamics [27], before crossing over to ordinary diffusion at longer times. For systems with long-range dipolar interactions, even smaller effective exponents in the range $\alpha\sim 0.6$$0.75$ have been reported [9].
The observed super-diffusive but sub-ballistic transport in our system therefore indicates dynamics beyond classical stochastic relaxation.
@@
Additional illustrations in the online only Appendix Fig. A1 and A2 extend this analysis to system with different charge densities and varying interaction conditions. The initial state is similar to popping a balloon filled with plasma in a vacuum. Increasing monopole plasma density or decreasing dipolar interactions causes the cloud to expand faster. The additional effect of dipolar interactions is evident in increased short-range correlations between positive and negative charges.
Qualitative comparison with expectations from $\mathbb{Z}_{2}$ lattice gauge theories suggests that the observed screening behavior, correlation decay, and absence of extended confining strings are consistent with a deconfined regime. We emphasize, however, that our measurements do not directly probe standard diagnostics such as Wilson loops or topological sector structure, and therefore do not constitute a complete identification of a $\mathbb{Z}_{2}$ spin liquid.
The observed dynamics provide direct access to the microscopic motion of spin-ice defects in a quantum-coherent regime of an artificial platform. While we do not yet have a quantitative theory for the measured transport exponents, the anomalous dynamics, already present in the nominally non-interacting regime, can be attributed to constraints of the Hilbert space, which generate effective entropic interactions between monopoles (see Supplementary Material for a coarse-grained derivation).
@@
Supplementary materials
@@
Movie of some of the configurations measured in the calibration of the energy scale of charge conservation enforcement. Initial state is the Dirac string shown in Fig. 2, configurations are measured after dynamics were run for $t_{eff}=100$.
Acknowledgements.
We thank Sherry Mayo for a careful reading of the manuscript and valuable comments. We thank Nilotpal Chakraborty for bringing the references about sub-diffusive scaling to our attention. P.S. acknowledges the funding from the Polish National Science Centre (NCN) Sonata Bis grant 2019/34/E/ST3/00405. Access to the D-Wave system was funded by the CSIRO Quantum Technologies Future Science Platform.
Author Contributions